OpenStudy (anonymous):

The formula for the area of a triangle is where b is the base of the triangle and h is the height of the triangle. What is the length of the base if the area is 22 and the height is 8 cm?

OpenStudy (zzr0ck3r):

1/2 (8) b = 22 can you solve for b?

OpenStudy (anonymous):

what do you mean

OpenStudy (zzr0ck3r):

(1/2)*8*x = 22 8x = 44 x = 11/2

OpenStudy (anonymous):

it has to be in decimal form and that cant be it

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

use a calculator 11/2 = 5.5

OpenStudy (anonymous):

wow my bad